OpenStudy (anonymous):

Three times the sum of twice a number and 3 is 33. Find the number.

OpenStudy (bibby):

2x+3=33 solve for x

OpenStudy (anonymous):

2x=33-3 2x=30 2x/2=30/2 x=15

OpenStudy (anonymous):

choices are The number is 24/5. b. The number is 4. c. The number is 6. d. The number is 5.